OpenStudy (anonymous):

Simplify: (y- (4/y))/(3+(6/y))

OpenStudy (anonymous):

\[\frac{y-\frac{4}{y}}{3+\frac{6}{y}}\] \[\frac{\frac{y^2-4}{y}}{\frac{3y+6}{y}}\] \[\frac{y^2-4}{\cancel y}*\frac{\cancel y}{3y+6}\] \[\frac{\cancel{(y+2)}(y-2)}{3\cancel{(y+2)}}\] \[\frac{y-2}{3}\]